# SCSA: Exploring the Synergistic Effects Between Spatial and Channel Attention

### v1

@@ Abstract
Channel and spatial attentions have respectively brought significant improvements in extracting feature dependencies and spatial structure relations for various downstream vision tasks. While their combination is more beneficial for leveraging their individual strengths, the synergy between channel and spatial attentions has not been fully explored, lacking in fully harness the synergistic potential of multi-semantic information for feature guidance and mitigation of semantic disparities. Our study attempts to reveal the synergistic relationship between spatial and channel attention at multiple semantic levels, proposing a novel Spatial and Channel Synergistic Attention module (SCSA). Our SCSA consists of two parts: the Shareable Multi-Semantic Spatial Attention (SMSA) and the Progressive Channel-wise Self-Attention (PCSA). SMSA integrates multi-semantic information and utilizes a progressive compression strategy to inject discriminative spatial priors into PCSA’s channel self-attention, effectively guiding channel recalibration. Additionally, the robust feature interactions based on the self-attention mechanism in PCSA further mitigate the disparities in multi-semantic information among different sub-features within SMSA. We conduct extensive experiments on seven benchmark datasets, including classification on ImageNet-1K, object detection on MSCOCO 2017, segmentation on ADE20K, and four other complex scene detection datasets. Our results demonstrate that our proposed SCSA not only surpasses the current state-of-the-art attention but also exhibits enhanced generalization capabilities across various task scenarios. The code and models are available at: https://github.com/HZAI-ZJNU/SCSA.
Index Terms:
Multi-semantic information, semantic disparity, spatial attention, channel attention, synergistic effect.
Figure 1: The accuracy of SCSA and various attention mechanisms is compared across multiple benchmark models on the ImageNet-1K validation set.
I Introduction
@@
Attention mechanisms, by enhancing representations of interest, facilitate the learning of more discriminative features and are widely used in redistributing channel relationships and spatial dependencies. Existing universal attention methods can be primarily categorized into three types: channel attention[1, 2, 3, 4, 5, 6], spatial attention[7, 8, 9, 10], and hybrid channel-spatial attention[11, 12, 13, 14, 15, 16, 17, 18]. Their focuses differ: channel attention concentrates on extracting features of objects, while spatial attention is tailored to augment spatial information. We notice that spatial information represents semantic feature objects at the pixel level. Local spatial information captures low-semantic objects, while global spatial information perceives high-semantic objects. Thus, different spatial information, through their semantic associations, reflects distinct feature objects which are distributed across various features. To this end, we hypothesize that spatial information can guide channel learning. To further investigate guidance methodologies, we first review the current mainstream attention mechanisms.
@@
CBAM[11] aggregates global spatial and channel information separately by chaining channel and spatial attention, but compressing all channel information leads to sharing across all spatial structures. This weakens the adaptability of spatial context to different feature maps. To overcome this, CPCA[19] introduces a channel-priority attention mechanism and depth-wise stripe convolutions, independently extracting spatial structures of each feature, significantly improving medical image segmentation. Furthermore, the EMA[17] module, based on grouped attention and cross-spatial multi-scale interactions, effectively integrates spatial information of both long and short-range dependencies but overlooks inter-group feature interactions. Although these hybrid attentions boost learning, they overlook the crucial guiding role of multi-semantic information in spatial-channel synergy.
@@
How can we construct a synergistic mechanism where spatial attention guides channel attention to enhance comprehensive learning, and channel attention modulates richer spatial-specific patterns from multi-semantic levels? Differing from the aforementioned methods, we explore the synergistic effects from three aspects: dimension decoupling, lightweight multi-semantic guidance, and semantic disparities mitigation, and propose a novel, plug-and-play Spatial and Channel Synergistic Attention(SCSA). Our SCSA is composed of a shareable Multi-Semantic Spatial Attention (SMSA) and a Progressive Channel Self-Attention (PCSA) linked sequentially. Our study initially employs multi-scale, depth-shared 1D convolutions to extract spatial information at various semantic levels from four independent sub-features. We utilize group normalization across four sub-features to hasten model convergence while avoiding the introduction of batch noise and the leakage of semantic information between sub-features. Subsequently, we input the SMSA-modulated feature maps into PCSA, incorporating progressive compression and channel-specific self-attention mechanisms (CSA). Our progressive compression strategy is designed to minimize computational complexity while preserving the spatial priors within SMSA, offering a practical trade-off. Moreover, our PCSA leverages an input-aware self-attention mechanism to effectively explore channel similarities, thereby mitigating semantic disparities among different sub-features in SMSA. We conducted extensive experiments across four visual tasks and seven benchmark datasets, illustrating the effectiveness of the multi-semantic synergy applied in our SCSA, and its superior generalization capability compared to other attention mechanisms. In summary, our contributions are as follows:
  • We propose an efficient SMSA that utilizes multi-scale depth-shared 1D convolutions to capture multi-semantic spatial information for each feature channel, effectively integrating global contextual dependencies and multi-semantic spatial priors.
  • Using SMSA, we develop feature-independent spatial structures and propose PCSA that calculates channel similarities and contributions guided by compressed spatial knowledge, mitigating semantic disparities in spatial structures.
  • We connected SMSA and PCSA in series to create the SCSA, exploring synergistic effects through dimension decoupling, lightweight multi-semantic guidance, and semantic disparities mitigation. Our experiments confirm its superiority over current state-of-the-art attention mechanisms in various visual tasks and complex scenarios.
@@
II Related Work
II-A Multi-Semantic Spatial Information
@@ II-B Attention Compression
Integrating attention mechanisms into various mainstream backbone or feature fusion networks enhances the model’s understanding of fine-grained features and accuracy in feature representation. However, it inevitably leads to increased memory usage and computational time. CA[15] performs unidirectional spatial compression along the height (H) and width (W) dimensions separately, preserving spatial structures in one direction while aggregating global spatial information in another, mitigating information loss from global compression. SA[16] and EMA[5] reshape features into sub-features, reducing attention computation and parameters. However, reshape operations constrained by GPU bandwidth can lead to costly data transfers, with considerable time spent on data rearrangement, impacting training and inference speeds. CPCA[19] uses stripe convolutions in independent channels to reduce parameters in large-kernel convolutions. Recent studies also apply dimension decomposition in MHSA, with RMT[30] applying MHSA separately across H and W dimensions to minimize computational costs.
Although some of the aforementioned attention methods have proven effective in specific domains, they still suffer performance degradation when generalized to more complex scenarios.
TABLE I: Comparison of our method, based on the UperNet model, with other attention mechanisms for semantic segmentation performance on the ADE20K benchmark.
Methods UperNet
Params(M) FLOPs(G) mIoU(%)
ResNet-50 64.10 1895 40.20
+ CBAM 66.62 1895 39.62
+ CPCA 65.94 1927 39.68
+ SE 66.62 1895 39.94
+ SA 64.10 1895 40.01
+ ECA 64.10 1895 40.46
+ FCA 66.61 1895 41.09
+ SCSA(Ours) 64.16 1895 41.14
ResNet-101 83.09 2051 42.74
+ CBAM 87.84 2051 41.65
+ ECA 83.09 2051 42.63
+ SE 87.84 2051 42.66
+ FCA 87.83 2051 43.22
+ SCSA(Ours) 83.22 2051 43.76
TABLE II: Comparison of our method, based on the Mask R-CNN, with other attention mechanisms for instance segmentation performance on MSCOCO val2017.
Methods Mask R-CNN
AP $AP_{50}$ $AP_{75}$ $AP_{S}$ $AP_{M}$ $AP_{L}$
ResNet50 34.8 55.9 36.9 16.4 37.4 50.2
+ CBAM 35.4 56.9 37.6 17.4 38.3 50.6
+ ECA 35.5 57.6 37.6 16.6 38.4 52.0
+ FCA 35.5 57.2 37.6 17.1 38.6 51.3
+ SE 35.7 57.3 38.1 17.7 38.6 50.9
+ SA 35.7 57.7 38.0 17.2 38.7 51.5
+ CA 35.8 57.5 38.2 16.9 38.5 51.7
+ SCSA 36.1 58.4 38.3 17.2 39.1 51.9
III Method
@@
In this section, we begin by discussing the SMSA module, which explores the benefits of lightweight multi-semantic information guidance. Next, we introduce the PCSA module, which utilizes a progressive compression strategy and channel-wise self-attention to mitigate semantic disparities. The synergistic effects of multi-semantic guidance and semantic disparities mitigation motivate us to propose SCSA module. The overall architecture is shown in Figure 2.
III-A Shared Multi-Semantic Spatial Attention
III-A1 Spatial and Channel Decomposition
@@
Decomposition techniques in neural network architectures significantly reduce the number of parameters and computational cost. Inspired by the structure of 1D sequences in NLP tasks, in our study, we decompose the given input ${X\in\mathbb{R}^{B\times C\times H\times W}}$ along the height and width dimensions. We apply global average pooling to each dimension, thereby creating two unidirectional 1D sequence structures: $X_{H}\in\mathbb{R}^{B\times C\times W}$ and $X_{W}\in\mathbb{R}^{B\times C\times H}$. To learn varying spatial distributions and contextual relationships, we partition the feature set into $K$ identically sized, independent sub-features, $X_{H}^{i}$ and $X_{W}^{i}$, with each sub-feature having a channel count of $\frac{C}{K}$. In this paper, we set the default value $K=4$. The process of decomposing into sub-features is presented as follows:
@@
III-A2 Lightweight Convolution Strategies Across Disjoint Sub-features
After partitioning the feature set into exclusive sub-features, we aim to efficiently capture distinct semantic spatial structures within each sub-feature. Inspired by extensive research[31, 25, 26] on reducing feature redundancy, which reveal that such redundancy is likely due to intense interactions among features, we also observe varied spatial structures among features. Based on these insight, and aiming to enrich semantic information, enhance semantic coherence, and minimize semantic gaps, we apply depth-wise 1D convolutions with kernel sizes of 3, 5, 7, and 9 in four sub-features. Furthermore, to address the limited receptive field caused by decomposing features into H and W dimensions and applying 1D convolutions separately, we use lightweight shared convolutions for alignment, implicitly modeling the dependency between the two dimensions by learning consistent features in both dimensions. The ablation details regarding them are provided in Table III. The implementation process for extracting multi-semantic spatial information is defined as follows:
@@ III-A2 Lightweight Convolution Strategies Across Disjoint Sub-features
To accurately compute spatial attention maps for each feature, we aggregate distinct semantic sub-features and use Group Normalization (GN) with $K$ groups for normalization. We opt for GN over the conventional Batch Normalization (BN) because our study finds that GN is superior in distinguishing semantic differences among sub-features. GN allows for the independent normalization of each sub-feature without introducing batch statistical noise, effectively mitigating semantic interference between sub-features and preventing attention dilution. This approach is validated by ablation studies shown in Table III. Finally, spatial attention is generated using a simple Sigmoid activation function, which activates and suppresses specific spatial regions. The computation of output features is as follows:
@@
$X_{s}=Attn_{H}\times Attn_{W}\times X$ (7)
@@
III-B Progressive Channel-wise Self-Attention
A prevalent approach to compute channel attention is through convolutional operations that explore dependencies among channels. There is limited theoretical research on how convolutional layers explicitly model these dependencies, aside from achieving it through backpropagation and gradient updates. Inspired by the significant advantages of the ViT[10] in utilizing MHSA for modeling similarities among different tokens in spatial attention calculations, we propose combining the MHSA concept with modulated spatial priors from SMSA to compute inter-channel similarities. Moreover, to preserve and utilize the multi-semantic spatial information extracted by SMSA, and to reduce the computational cost of MHSA, we employ a progressive compression method, which reflects the guidance in our synergistic effects. Compared to traditional convolutional operations, PCSA exhibits stronger input perception capabilities and effectively utilizes the spatial priors provided by SMSA to deepen learning. The implementation details of our PCSA are as follows:
@@ III-B Progressive Channel-wise Self-Attention
$X_{c}=X_{s}\times\sigma(Pool_{(H^{\prime},W^{\prime})}^{(H^{\prime},W^{\prime})\rightarrow(1,1)}(X_{attn}))$ (12)
$Pool_{(7,7)}^{(H,W)\rightarrow(H^{\prime},W^{\prime})}(\cdot)$ denotes a pooling operation with a kernel size of 7x7 that rescales the resolution from $(H,W)$ to $(H^{\prime},W^{\prime})$. $F_{proj}(\cdot)$ represents the mapping function that generates the query, key, and value.
It’s important to note that, unlike the MHSA in the ViTs where $Q,K,V\in\mathbb{R}^{B\times N\times C}$ with $N=HW$, in our PCSA’s CA-MHSA, self-attention is computed along the channel dimension, with $Q,K,V\in\mathbb{R}^{B\times C\times N}$.
III-C Synergistic Effects
The synergistic spatial and channel attention mechanisms aim to complement each other. In our work, we propose a novel concept of guiding channel attention learning through spatial attention. Drawing from the connection methods of CBAM[11] and CPCA[19], we employ a simple serial connection to integrate our SMSA and PSCA modules. Our innovation lies in the meticulously designed spatial and channel attentions: spatial attention extracts multi-semantic spatial information from each feature, providing precise spatial priors for channel attention computation; channel attention refines the semantic understanding of local sub-feature $X^{i}$ by leveraging the overall feature map $X$, mitigating semantic disparities caused by multi-scale convolution in SMSA. Additionally, unlike previous approaches[1, 11, 15, 12], we do not employ channel compression, effectively preventing the loss of crucial features. Ultimately, our constructed SCSA is as follows:

### v2

@@ Abstract
Channel and spatial attentions have respectively brought significant improvements in extracting feature dependencies and spatial structure relations for various downstream vision tasks. The combined use of both channel and spatial attentions is widely considered beneficial for further performance improvement; however, the synergistic effects between channel and spatial attentions, especially in terms of spatial guidance and mitigating semantic disparities, have not yet been thoroughly studied. This motivates us to propose a novel Spatial and Channel Synergistic Attention module (SCSA), entailing our investigation toward the synergistic relationship between spatial and channel attentions at multiple semantic levels. Our SCSA consists of two parts: the Shareable Multi-Semantic Spatial Attention (SMSA) and the Progressive Channel-wise Self-Attention (PCSA). SMSA integrates multi-semantic information and utilizes a progressive compression strategy to inject discriminative spatial priors into PCSA’s channel self-attention, effectively guiding channel recalibration. Additionally, the robust feature interactions based on the Channel-wise single-head self-attention mechanism in PCSA further mitigate the disparities in multi-semantic information among different sub-features within SMSA. We conduct extensive experiments on seven benchmark datasets, including classification on ImageNet-1K, object detection on MSCOCO, segmentation on ADE20K, and four other complex scene detection datasets. Our results demonstrate that our proposed SCSA not only surpasses the current plug-and-play state-of-the-art attention but also exhibits enhanced generalization capabilities across various task scenarios. The code and models are available at: https://github.com/HZAI-ZJNU/SCSA .
keywords:
[…]
1 Introduction
@@
Attention mechanisms, by enhancing representations of interest, facilitate the learning of more discriminative features and are widely used in redistributing channel relationships and spatial dependencies. Existing plug-and-play attention methods can be primarily categorized into three types: channel attention [23, 27, 62, 55, 31, 41], spatial attention [22, 56, 42, 9], and hybrid channel-spatial attention [57, 40, 38, 14, 20, 65, 39, 61]. Their focuses differ: channel attention focus on enhancing the extraction of key object features by adaptively weighting the importance of different channels, while spatial attention is tailored to augment critical spatial information. Spatial information represents semantic feature objects at the pixel level. Local spatial information captures low-semantic objects, such as fine details and textures, while global spatial information perceives high-semantic objects, such as overall shape.
Figure 1: Visualization of several feature maps. Different box or circle colors highlight inherent spatial semantic disparities across specific parts in various feature maps.
@@
In models based on convolutional neural network (CNN) architectures, deep convolutional operators are commonly employed for feature extraction. The gradients generated by these operators can flow and propagate across different feature channels, facilitating the update of convolutional weights and effectively representing image features. However, numerous studies have shown that relying solely on the convolutional branch for feature extraction can lead to distorted gradient flows [54], resulting in the loss of critical information or redundancy of similar features [4, 17, 66, 35]. To address this, several methods based on excitation and suppression mechanisms [23] have been proposed, which focus feature learning on the most critical features for different tasks, thereby enhancing the model’s representational capacity. In particular, CBAM [57] aggregates global spatial and channel information separately by chaining channel and spatial attention, but compressing all channel information leads to sharing across all spatial structures. This weakens the adaptability of spatial context to different feature maps. To overcome this, CPCA [24] introduces a channel-priority attention mechanism and depth-wise stripe convolutions, independently extracting spatial structures of each feature, significantly improving medical image segmentation. Furthermore, the EMA [39] module, based on grouped attention and cross-spatial multi-scale interactions, effectively integrates spatial information of both long and short-range dependencies but overlooks inter-group feature interactions.
Although these hybrid attention mechanisms enhance representation learning, they overlook the inherent multi-semantic information across spatial and channel dimensions, as well as the interaction and disparity mitigation of multi-semantic features, which are crucial for fine-grained tasks such as detection and segmentation, thereby limiting the plug-and-play capability of these methods. As shown in fig. 1, we analyze several feature maps of the image and observe that distinct spatial regions exhibit inherent semantic disparities and similarities, arising from the feature selectivity of different channels.
Based on this insight, this raises the question of whether it is possible to leverage the inherent spatial semantic disparities across different feature channels to guide the learning of important features? Furthermore, given the presence of semantic disparities, how can we mitigate these multi-semantic differences and promote better fusion of multi-semantic information?
@@
Differing from the aforementioned methods, we explore solutions to the above issues from following three aspects: dimension decoupling, lightweight multi-semantic guidance, and semantic disparities mitigation, and propose a novel, plug-and-play Spatial and Channel Synergistic Attention (SCSA). Our SCSA is composed of a shareable Multi-Semantic Spatial Attention (SMSA) and a Progressive Channel-wise Self-Attention (PCSA) linked sequentially. Our study initially employs multi-scale, depth-shared 1D convolutions to extract spatial information at various semantic levels from four independent sub-features. We utilize Group Normalization [58] across four sub-features to hasten model convergence while avoiding the introduction of batch noise and the interference of semantic information between different sub-features. Subsequently, we input the SMSA-modulated feature maps into PCSA, incorporating progressive compression and channel-specific single-head self-attention mechanisms. Our progressive compression strategy is designed to minimize computational complexity while preserving the spatial priors within SMSA, offering a practical trade-off. Moreover, our PCSA leverages an input-aware single-head self-attention mechanism to effectively explore channel similarities, thereby mitigating semantic disparities among different sub-features in SMSA and promoting information fusion. We conducted extensive experiments across four visual tasks and seven benchmark datasets, illustrating the effectiveness of the multi-semantic synergy applied in our SCSA. In summary, our contributions are as follows:
  • We identify two key limitations in existing plug-and-play attention mechanisms: 1) insufficient utilization of inherent multi-semantic spatial information to guide the extraction of key features along spatial and channel dimensions, and 2) inadequate handling of semantic disparities and interactions caused by multi-semantic information across feature maps.
  • We propose the Spatial and Channel Synergistic Attention (SCSA), comprising the SMSA and PCSA modules. SMSA utilizes multi-scale depth-shared 1D convolutions to capture multi-semantic spatial information, enhancing both local and global feature representations. PCSA employs input-aware self-attention to refine channel features, effectively mitigating semantic disparities and ensuring robust feature integration across channels.
  • Our proposed method outperforms other state-of-the-art plug-and-play attention mechanisms on multiple benchmarks, including ImageNet-1K for classification, MSCOCO for object detection, and ADE20K for segmentation, and demonstrates strong generalization capability across various complex scenarios, such as low-light and small-object benchmarks.
@@
2 Related Work
2.1 Multi-Semantic Spatial Information
@@ 2.2 Attention Decomposition
Integrating attention mechanisms into mainstream backbones or feature fusion networks enhances fine-grained feature understanding and improves feature representation accuracy. However, it inevitably leads to increased memory usage and computational time. CA [20] and ELA [61] perform unidirectional spatial compression along the height (H) and width (W) dimensions separately, preserving spatial structures in one direction while aggregating global spatial information in another, mitigating information loss from global compression. SA [65] and EMA [39] reshape features into sub-features, reducing attention computation and parameters. However, the reshape operations they used in the high-dimensional B (batch size) and C (number of channels) constrained by GPU bandwidth can lead to expensive data transfers, significantly impacting inference speeds. CPCA [24] uses stripe convolutions in independent channels to reduce parameters in large-kernel convolutions. Recent studies also apply dimension decomposition in MHSA, with RMT [12] applying MHSA separately across H and W dimensions to minimize computational costs.
In this study, we build upon the concept of attention decomposition and propose a lightweight guidance module that integrates multi-semantic spatial information. Additionally, we design a multi-semantic discrepancy mitigation module based on a progressive channel-wise single-head self-attention mechanism, aiming to explore a more optimized synergistic relationship between the spatial and channel dimensions.
3 Method
@@
In this section, we begin by discussing the SMSA module, which explores the benefits of lightweight multi-semantic information guidance. Next, we introduce the PCSA module, which utilizes a progressive compression strategy and channel-wise self-attention to mitigate semantic disparities. The synergistic effects of multi-semantic guidance and semantic disparities mitigation motivate us to propose SCSA module. The overall architecture is shown in fig. 2.
3.1 Shared Multi-Semantic Spatial Attention
Spatial and Channel Decomposition.
@@ 3.1 Shared Multi-Semantic Spatial Attention
Decomposition techniques in neural network architectures substantially reduce the parameter count and computational overhead. Inspired by the structure of 1D sequences in Transformer [52], in our work, we decompose the given input ${X\in\mathbb{R}^{B\times C\times H\times W}}$ along the height and width dimensions. We apply global average pooling to each dimension, thereby creating two unidirectional 1D sequence structures: $X_{H}\in\mathbb{R}^{B\times C\times W}$ and $X_{W}\in\mathbb{R}^{B\times C\times H}$. To learn varying spatial distributions and contextual relationships, we partition the feature set into $K$ identically sized, independent sub-features, $X_{H}^{i}$ and $X_{W}^{i}$, with each sub-feature having a channel count of $\frac{C}{K}$. In this paper, we set the default value $K=4$. The process of decomposing into sub-features is presented as follows:
@@
Lightweight Convolution Strategies Across Disjoint Sub-features.
After partitioning the feature maps into exclusive sub-features, we aim to efficiently capture distinct semantic spatial structures within each sub-feature. Inspired by extensive research on reducing feature redundancy [35, 17, 4], which reveal that such redundancy is likely due to intense interactions among features, we also observe varied spatial structures among features, as illustrated in fig. 1. Based on these insights and aiming to enrich semantic information, enhance semantic coherence, and minimize semantic gaps, we apply depth-wise 1D convolutions with kernel sizes of 3, 5, 7, and 9 to four sub-features. Furthermore, to address the limited receptive field caused by decomposing features into H and W dimensions and applying 1D convolutions separately, we use lightweight shared convolutions for alignment, implicitly modeling the dependency between the two dimensions by learning consistent features across both. The ablation details regarding them are provided in table 1. The implementation process for extracting multi-semantic spatial information is defined as follows:
@@
After decomposing the independent sub-features and capturing the spatial information of different semantics, we need to construct the spatial attention map. Specifically, we concat distinct semantic sub-features and use Group Normalization (GN) [58] with $K$ groups for normalization. We opt for GN over the common Batch Normalization (BN) [26] because our study finds that GN is superior in distinguishing semantic differences among sub-features. GN allows for the independent normalization of each sub-feature without introducing batch statistical noise, effectively mitigating semantic interference between sub-features and preventing attention dilution. This design is validated by ablation studies shown in table 1. Finally, spatial attention is generated using a simple Sigmoid activation function, which activates and suppresses specific spatial regions. The computation of output features is as follows:
@@
$SMSA(X)=X_{s}=Attn_{H}\times Attn_{W}\times X$ (7)
@@
3.2 Progressive Channel-wise Self-Attention
A prevalent approach to compute channel attention is through convolutional operations that explore dependencies among channels [23, 55]. The use of convolution to model the similarities between features is somewhat non-intuitive and makes it difficult to effectively measure the similarity across different channels. Inspired by the significant advantages of the ViT [9] in utilizing MHSA for modeling similarities among different tokens in spatial, we propose combining the Single-Head Self-Attention (SHSA) with modulated spatial priors from SMSA to compute inter-channel similarities. Moreover, to preserve and utilize the multi-semantic spatial information extracted by SMSA, and to reduce the computational cost of SHSA, we employ a progressive compression method based on average pooling, which serves as the guidance in our synergistic effects. Compared with modeling channel dependencies using common convolutional operations, PCSA exhibits stronger input perception capabilities and effectively leverages the spatial priors provided by SMSA to deepen learning. The implementation details of our PCSA are as follows:
@@ 3.2 Progressive Channel-wise Self-Attention
$PCSA(X_{s})=X_{c}=X_{s}\times\sigma(Pool_{(H^{\prime},W^{\prime})}^{(H^{\prime},W^{\prime})\rightarrow(1,1)}(X_{attn}))$ (12)
$Pool_{(k,k)}^{(H,W)\rightarrow(H^{\prime},W^{\prime})}(\cdot)$ denotes a pooling operation with a kernel size of $k\times k$ that rescales the resolution from $(H,W)$ to $(H^{\prime},W^{\prime})$. $F_{proj}(\cdot)$ represents the linear projection that generates the query, key, and value.
It’s important to note that, unlike the MHSA in the ViTs where $Q,K,V\in\mathbb{R}^{B\times N\times C}$ with $N=HW$, in our PCSA’s CA-SHSA, self-attention is computed along the channel dimension, with $Q,K,V\in\mathbb{R}^{B\times C\times N}$. Additionally, to fully interact with the different sub-features decomposed in SMSA, We select to implement a simpler single-head self-attention mechanism instead of the combination of multi-head self-attention with channel shuffling [66].
3.3 Synergistic Effects
The synergistic spatial and channel attention mechanisms aim to complement each other. In our work, we propose a novel concept of guiding channel attention learning through spatial attention. Drawing inspiration from the connection between CBAM [57] and CPCA[24], we employ a similar serial structure to integrate our SMSA and PSCA modules, forming the Spatial and Channel Synergistic Attention (SCSA). The difference is that the spatial attention SMSA is applied first, followed by the channel attention PSCA. The former extracts multi-semantic spatial information from each feature, providing precise spatial priors for the latter; the latter refines the semantic understanding of the local sub-feature $X^{i}$ by leveraging the overall feature map $X$, thereby mitigating the semantic disparities caused by the multi-scale convolutions in the former. Additionally, unlike previous approaches [23, 57, 40, 20], we do not employ channel compression, effectively preventing the loss of crucial features. Ultimately, our constructed SCSA is as follows: